# PAPER: Privacy-Preserving ResNet Models using Low-Degree Polynomial Approximations and Structural Optimizations on Leveled FHE

### v1

@@ Abstract
Recent work has made non-interactive privacy-preserving inference more practical by running deep Convolution Neural Network (CNN) with Fully Homomorphic Encryption (FHE). However, these methods remain limited by their reliance on bootstrapping, a costly FHE operation applied across multiple layers, severely slowing inference. They also depend on high-degree polynomial approximations of non-linear activations, which increase multiplicative depth and reduce accuracy by 2–5% compared to plaintext ReLU models. In this work, we focus on ResNets, a widely adopted benchmark architecture in privacy-preserving inference, and close the accuracy gap between their FHE-based non-interactive models and plaintext counterparts, while also achieving faster inference than existing methods. We use a quadratic polynomial approximation of ReLU, which achieves the theoretical minimum multiplicative depth for non-linear activations, along with a penalty-based training strategy. We further introduce structural optimizations such as node fusing, weight redistribution, and tower reuse. These optimizations reduce the required FHE levels in CNNs by nearly a factor of five compared to prior work, allowing us to run ResNet models under leveled FHE without bootstrapping. To further accelerate inference and recover accuracy typically lost with polynomial approximations, we introduce parameter clustering along with a joint strategy of data encoding layout and ensemble techniques. Experiments with ResNet-18, ResNet-20, and ResNet-32 on CIFAR-10 and CIFAR-100 show that our approach achieves up to $4\times$ faster private inference than prior work with comparable accuracy to plaintext ReLU models.
@@ 1 Introduction
Machine Learning as a Service (MLaaS) is increasingly adopted across industries as it provides access to powerful models without the need for in-house development or maintenance (Ribeiro et al., 2015). However, it raises serious privacy concerns since models are often trained on proprietary or sensitive data (Anwar et al., 2018; Özbayoglu et al., 2020), and inference requires clients to share private information (e.g., medical or financial records) with external service providers. Privacy-Preserving Machine Learning (PPML) addresses these risks using cryptographic methods, mainly secure Multi-Party Computation (MPC) (Zhou et al., 2024) and Fully Homomorphic Encryption (FHE) (Podschwadt et al., 2022), which protect both client data and provider models. Broadly, PPML methods fall into interactive and non-interactive categories (Lee et al., 2022). Interactive PPML, based on MPC, requires the client and the service provider to jointly perform inference (Liu et al., 2017; Juvekar et al., 2018; Mishra et al., 2020; Lou et al., 2021; Huang et al., 2022; Diaa et al., 2024). It keeps computational costs relatively low but requires multiple communication rounds, leading to higher communication and bandwidth requirements that limit applicability. Non-interactive PPML, typically using FHE, adopts a fire-and-forget paradigm (Gilad-Bachrach et al., 2016; Brutzkus et al., 2019; Lou & Jiang, 2021; Lee et al., 2022; Sarkar et al., 2023; Ao & Boddeti, 2024), where the client encrypts inputs, the provider computes on ciphertexts, and the client decrypts the output. This reduces communication to a single round at the expense of heavy server-side computation. Despite the costs, non-interactive PPML is well-suited to third-party services as it eliminates client involvement during inference, supports low-resource devices, and works in low-bandwidth settings. Motivated by these advantages, this work focuses on non-interactive PPML.
@@
A key challenge in non-interactive PPML is implementing activation functions with cryptographic primitives. Activations are non-polynomial, while FHE supports only polynomial operations like addition and multiplication. Prior work addresses this by approximating activations with single polynomials (Diaa et al., 2024) or piecewise polynomials (Liu et al., 2017). These approximations can be introduced before or after training. In post-training approximation, models are trained with standard activations (e.g., ReLU) and later replaced with high-precision polynomials. For example, MPCNN (Lee et al., 2022) uses minimax polynomials of orders $\{15,15,27\}$. In pre-training approximation, ReLU is replaced with a polynomial before training, allowing lower-degree approximations that reduce multiplicative depth and speed up inference. For example, shallow CNNs like CryptoNets (Gilad-Bachrach et al., 2016) and LoLa (Brutzkus et al., 2019), similar to LeNet-5 (LeCun et al., 1998), achieve over 98% accuracy on MNIST with degree-2 polynomials. However, Garimella et al. (2021) showed that training larger models (beyond AlexNet (Krizhevsky et al., 2012) or VGG-11 (Simonyan & Zisserman, 2015)) with such low-degree polynomials causes large approximation errors that destabilize training. To date, PILLAR (Diaa et al., 2024), which uses a degree-4 polynomial, is the lowest-degree approximation that generalizes to deep neural networks.
@@
Implementing non-interactive PPML becomes increasingly difficult as model depth increases. While bootstrapping enables evaluation of arbitrarily deep models by refreshing ciphertexts, it is an expensive operation in FHE and remains the main scalability bottleneck. To avoid this cost, prior work has relied on Leveled Fully Homomorphic Encryption (LFHE) (see §A for details). LFHE eliminates bootstrapping and allows faster private inference, but its computation is bounded by a fixed multiplicative depth determined by encryption parameters. As a result, the size and complexity of neural networks that can be evaluated under LFHE are strictly limited. The deepest CNNs shown to run with LFHE alone are AlexNet and VGG-11, as demonstrated by Garimella et al. (2021). Larger models such as ResNet-20 or ResNet-32 have so far required bootstrapping. For example, AutoFHE executes ResNet-20 inference with at least five bootstrapping layers (Ao & Boddeti, 2024), which leads to heavy computational costs. In this work, we focus on ResNet (He et al., 2016), as it is widely adopted in state-of-the-art evaluations and is well-suited for efficient and low-latency deployment.
@@
Contributions: This work advances non-interactive PPML by enabling, for the first time, large CNNs like ResNet-18, ResNet-20, and ResNet-32 to run entirely under LFHE without bootstrapping, achieving accuracy comparable to plaintext ReLU models, while outperforming prior work in both accuracy and inference time. Our contributions are threefold:
@@
  • To the best of our knowledge, this is the first work to train large CNNs, like ResNet-18, ResNet-20, and ResNet-32 effectively with degree-2 polynomial activations. We introduce a novel penalty function that ensures stable training and accuracy on par with ReLU-based models. This achieves the theoretical minimum multiplicative depth of one for non-linear activations, whereas the lowest previously known stable alternative, PILLAR, required depth three (see §2).
@@
  • We introduce three novel structural optimization techniques that reduce the overall multiplicative depth of a model while preserving functional equivalence: (i) Node Fusing, which merges consecutive operations such as convolution and batch normalization by folding normalization parameters into convolution weights, eliminating separate multiplications; (ii) Weight Redistribution, which adjusts parameters so that coefficients of highest-order terms in polynomial activations and batch normalization, as well as divisors in pooling layers, normalize to one, removing redundant multiplications by constants; and (iii) Tower Reuse, which allows multiple multiplications within the same FHE level before rescaling. In the CKKS scheme used in this work (Cheon et al., 2017), ciphertexts carry a scale factor that grows with multiplications. Rescaling keeps ciphertext values within range but consumes a level. Since each level requires a modulus and homomorphic operations apply to all moduli, fewer rescaling steps reduce the number of moduli, leading to faster homomorphic operations. Moreover, they permit using larger moduli for higher accuracy (see §3).
  • We propose an encoding strategy that utilizes unused ciphertext slots in FHE to pack multiple model instances, enabling simultaneous ensemble inference to improve accuracy at some additional inference cost. To offset this overhead, we introduce a parameter clustering method for convolutions that recovers much of the lost speed. To the best of our knowledge, this is the first work to exploit unallocated ciphertext slots for ensemble inference in FHE (see §4).
@@
We evaluate our method on CIFAR-10 and CIFAR-100 datasets (Krizhevsky & Hinton, 2009), and show that it achieves up to $4\times$ faster private inference than prior work while achieving accuracy comparable to plaintext ReLU models (see §5).
2 Model Training with Polynomial Approximation
@@ 2.1 Problem Setting
Dataset. We consider a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{M}$ for multi-class classification, where $x_{i}\in\mathbb{R}^{n}$ is the feature vector and $y_{i}\in\{1,\dots,K\}$ is the class label. $\mathcal{D}$ has $M$ samples and $K$ classes.
ReLU Network. Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{K}$ be an $L$-layer feed-forward network with ReLU activations, parameterized by weight matrices $W^{(1)},\dots,W^{(L)}$, where each $W^{(l)}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ and $n_{l}$ is the number of neurons in layer $l$. The network output is given by $f(x)=h^{(L)}(x)$, where $h^{(0)}(x)=x$ and $h^{(l)}(x)=ReLU(W^{(l)}h^{(l-1)}(x))$ for all $1\leq l\leq L$. The activation function $ReLU(z)$ is applied coordinate-wise and defined as $ReLU(z)=\max(0,z)$.
Polynomial Network. We define $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{K}$ with the same architecture as $f$, but replace ReLU with a degree-$d$ polynomial $p_{d}(z)=\sum_{k=0}^{d}a_{k}z^{k}$ that approximates ReLU over a bounded interval [c,c][-c,c]. The polynomial coefficients $\{a_{k}\}_{k=0}^{d}$ are chosen so that the maximum approximation error satisfies: $\max_{z\in[-c,c]}|ReLU(z)-p_{d}(z)|\leq\varepsilon$, where $\varepsilon>0$ is a small constant. The network output is given by $g(x)=h^{(L)}_{p_{d}}(x)$, where $h^{(0)}_{p_{d}}(x)=x$ and $h^{(l)}_{p_{d}}(x)=p_{d}(W^{(l)}h^{(l-1)}_{p_{d}}(x))$ for all $1\leq l\leq L$.
Challenges. Using polynomial networks ensures FHE compatibility but introduces training challenges such as escaping activations and coefficient truncation (see §B for more details). Escaping activations occur when intermediate outputs drift outside the interval [c,c][-c,c], leading to large approximation errors. Coefficient truncation occurs when polynomial coefficients are stored with limited fixed-point precision $b$, which changes the polynomial’s shape and causes deviation from ReLU even inside the approximation interval. These issues require careful selection of $d$, [c,c][-c,c], and $b$, together with regularization during training to keep activations within the valid approximation range.

### v2

# PAPER: Privacy-Preserving Convolutional Neural Networks using Low-Degree Polynomial Approximations and Structural Optimizations on Leveled FHE

@@ Abstract
Recent work using Fully Homomorphic Encryption (FHE) has made non-interactive privacy-preserving inference of deep Convolutional Neural Networks (CNN) possible. However, the performance of these methods remain limited by their heavy reliance on bootstrapping, a costly FHE operation applied across multiple layers, severely slowing inference. Moreover, they depend on high-degree polynomial approximations of non-linear activations, which increase multiplicative depth and reduce accuracy by 2–5% compared to plaintext ReLU models. In this work, we close the accuracy gap between FHE-based non-interactive CNNs and their plaintext counterparts, while also achieving faster inference than existing methods. We propose a quadratic polynomial approximation of ReLU, which achieves the theoretical minimum multiplicative depth for non-linear activations, together with a penalty-based training strategy. We further introduce structural optimizations that reduce the required FHE levels in CNNs by a factor of five compared to prior work, allowing us to run deep CNN models under leveled FHE without bootstrapping. To further accelerate inference and recover accuracy typically lost with polynomial approximations, we introduce parameter clustering along with a joint strategy of data layout and ensemble techniques. Experiments with VGG and ResNet models on CIFAR and Tiny-ImageNet datasets show that our approach achieves up to $4\times$ faster private inference than prior work, with accuracy comparable to plaintext ReLU models.
@@ 1 Introduction
Machine Learning as a Service (MLaaS) is increasingly adopted across industries because it provides access to powerful models without requiring in-house development or maintenance [30]. However, it raises serious privacy concerns since models are often trained on proprietary or sensitive data [4, 28], and inference requires clients to share private information (e.g., medical or financial records) with external service providers. Privacy-Preserving Machine Learning (PPML) addresses these risks using predominantly cryptographic methods, mainly secure Multi-Party Computation (MPC) [34] and Fully Homomorphic Encryption (FHE) [29], which protect both client data and provider models. Broadly, PPML methods fall into interactive and non-interactive categories [20]. Interactive PPML, based on MPC, requires the client and the service provider to jointly perform inference [21, 16, 27, 24, 14, 11]. It keeps computational costs relatively low but requires multiple communication rounds, which leads to high communication and bandwidth requirements. Non-interactive PPML, typically using FHE, adopts a fire-and-forget paradigm [13, 6, 23, 20, 31, 5], where the client encrypts inputs, the provider computes on ciphertexts, and the client decrypts the output. This reduces communication to a single round at the expense of heavy server-side computation. Despite the costs, non-interactive PPML is well-suited to third-party services because it eliminates client involvement during inference, supports clients with low-resource devices, and works in low-bandwidth settings. Motivated by these advantages, this work focuses on non-interactive PPML.
@@
A key challenge in non-interactive PPML is implementing activation functions with cryptographic primitives. Activations are non-polynomial, and FHE natively supports only polynomial operations such as addition and multiplication. Existing work addresses this by approximating activations with single polynomials [11] or piecewise polynomials [21]. These approximations can be introduced before or after training. In post-training approximation, models are trained with standard activations (e.g., ReLU) and later replaced with high-precision polynomials. For example, MPCNN [20] uses minimax polynomials of orders $\{15,15,27\}$. In pre-training approximation, ReLU is replaced with a polynomial before training, which allows lower-degree approximations that reduce multiplicative depth and speed up inference. For example, shallow CNNs like CryptoNets [13] and LoLa [6], similar to LeNet-5 [19], achieve over 98% accuracy on MNIST with degree-2 polynomials. However, prior work [12] showed that training larger models (beyond AlexNet [18] or VGG-11 [33]) with such low-degree polynomials causes large approximation errors that destabilize training. To date, PILLAR [11], which uses a degree-4 polynomial, is the lowest-degree approximation that generalizes to deep neural networks.
@@
Implementing non-interactive PPML becomes increasingly difficult with deeper models. While bootstrapping enables evaluation of arbitrarily deep models, it is an expensive operation in FHE and remains the main scalability bottleneck. To avoid this cost, the literature has relied on Leveled Fully Homomorphic Encryption (LFHE). LFHE eliminates bootstrapping and allows faster private inference, but its computation is bounded by a fixed multiplicative depth determined by encryption parameters. As a result, the size and complexity of neural networks that can be evaluated under LFHE are strictly limited. The deepest CNNs shown to run with LFHE alone are AlexNet and VGG-11 [12]. Larger models have so far required bootstrapping. Consequently, modern deep CNNs have largely been regarded as incompatible with LFHE, and bootstrapping has become the default mechanism to address this limitation despite its high computational cost. For example, AutoFHE executes ResNet-32 inference with 8 or 19 bootstrapping layers in its lower- or higher-accuracy configuration, respectively [5], resulting in heavy computational costs.
This paper revisits the prevailing assumption that modern deep CNNs are incompatible with LFHE. We argue that heavy reliance on bootstrapping has obscured a large, underexplored optimization space. In FHE-based PPML, multiplicative depth directly determines the encryption parameters and the frequency of bootstrapping operations needed during evaluation, both of which have a dominant impact on inference time. Crucially, the multiplicative depth of CNNs is not an immutable property of the model architecture itself, but emerges from how its computation is realized. The network structure, degree of polynomial activations, linear transformations, data layout, function implementations, and placement of rescaling operations collectively govern depth. Careful analysis and optimization of these components and their interactions can substantially reduce multiplicative depth without changing model functionality. By aggressively minimizing multiplicative depth at every level of the model, rather than compensating for inefficiencies with bootstrapping, we show that deep CNNs can be evaluated efficiently under LFHE.
Guided by this philosophy, our goal is to design the fastest non-interactive PPML framework that does not compromise accuracy, under LFHE. We achieve this by combining four techniques: (i) low-degree polynomial activations that attain the theoretical minimum multiplicative depth for non-linear activation functions, (ii) a suite of structural optimizations that remove redundant operations, (iii) an ensemble-based strategy that recovers accuracy to near-ReLU levels, and (iv) a clustering technique to speed up the computation of convolutional layers. Importantly, while this work focuses on LFHE, the proposed techniques are not limited to it. They can be applied with FHE to run very deep models while substantially reducing the number of required bootstrapping operations.
@@
Contributions: This work advances non-interactive PPML by enabling, for the first time, deep CNNs (e.g. ResNet-32) to run entirely under LFHE without bootstrapping, achieving accuracy comparable to plaintext ReLU models while outperforming prior work in both accuracy and inference time. Our contributions are fourfold:
@@
  • To the best of our knowledge, this is the first work to train deep CNNs effectively with degree-2 polynomial activations. We introduce a novel penalty function that ensures stable training and delivers accuracy on par with plaintext ReLU-based models. This achieves the theoretical minimum multiplicative depth of one for non-linear activations, whereas the lowest previously known stable alternative, PILLAR, requires a multiplicative depth of three (see §3).
@@
  • We propose three structural optimization techniques that reduce multiplicative depth of a model while preserving functional equivalence: (i) Node Fusing, which merges consecutive layers by folding normalization parameters into neighboring nodes, effectively eliminating the normalization layer; (ii) Weight Redistribution, which adjusts parameters so that coefficients of highest-order terms in polynomial activations and batch normalization, as well as divisors in pooling layers, normalize to one, removing redundant multiplications by constants; and (iii) Tower Reuse, which allows multiple multiplications within the same FHE level before rescaling. In the CKKS scheme used in this work [8], ciphertexts carry a scaling factor that grows with multiplications. Rescaling keeps ciphertext values within range but consumes a level. Since each level is represented by a modulus and homomorphic operations apply to all moduli, requiring fewer rescaling steps reduce the number of moduli, leading to faster homomorphic operations (see §4).
  • We introduce a novel encoding strategy that utilizes unused ciphertext slots in FHE to pack multiple CNN model instances, enabling simultaneous ensemble inference within a single homomorphic evaluation. This approach enables the use of multiple independently trained models to recover accuracy lost to low-degree polynomial approximations through careful reuse of ciphertext–plaintext structures, albeit with additional inference overhead (see §5).
  • To offset the additional private inference overhead from the choice of data layout that enables ensemble inference, we introduce a parameter clustering technique for convolutional layers that groups similar parameters to reduce redundant plaintext encodings and homomorphic computation. This technique recovers much of the lost inference speed while preserving model accuracy (see §5).
@@
We evaluate our method on CIFAR [17] and Tiny-ImageNet [10] datasets, and show that it achieves up to $4\times$ faster private inference than prior work while achieving accuracy comparable to plaintext ReLU models (see §6).
2 Background
2.1 Leveled Fully Homomorphic Encryption
Homomorphic Encryption allows meaningful computations to be performed directly on encrypted data. In this work, we employ Residue Number System (RNS) variant [7] of the Cheon-Kim-Kim-Song (CKKS) [8] encryption scheme, which supports approximate arithmetic over encrypted complex numbers. CKKS is an FHE scheme based on the Ring-Learning With Errors (RLWE) problem [25]. The scheme operates across three spaces: message $\mathcal{M}=\mathbb{C}$, plaintext $\mathcal{P}=R_{Q}=\mathbb{Z}_{Q}[x]/(x^{N}+1)$, and ciphertext $\mathcal{C}=R_{Q}\times R_{Q}$. Encoding maps a vector of $N/2$ numbers $\in\mathcal{M}$ into a degree-$N$ polynomial $\in\mathcal{P}$, and encryption transforms a plaintext $\in\mathcal{P}$ into a ciphertext $\in\mathcal{C}$. In RNS-CKKS, the ciphertext modulus $Q$ is decomposed into $\mathcal{L}+1$ smaller primes $Q=\prod_{i=0}^{\mathcal{L}}{q_{i}}$, where $\mathcal{L}$ denotes the level. A polynomial in $R_{Q}$ is represented as $\mathcal{L}+1$ polynomials in $R_{q_{i}}\forall i\in[0,\mathcal{L}]\cap\mathbb{Z}$, which allows computations using native 64-bit integer arithmetic.
CKKS supports homomorphic additions, multiplications, and rotations. Additions and multiplications act element-wise on the encoded vector, akin to SIMD operations. Rotations rotate the vector of $N/2$ encoded messages by a step $\in\mathbb{Z}_{N/2}$. During encoding, coefficients are scaled by a factor $\Delta$, rounded to the nearest integers, and reduced modulo $Q$. Multiplying two encoded messages scales the result to $\Delta^{2}$, requiring rescaling to restore the scale to $\Delta$. Rescaling removes one RNS modulus, reducing the level by 1. When the level reaches zero, bootstrapping can reset the ciphertext to a higher level, enabling unbounded computation. However, it is computationally expensive and avoided when possible. FHE without bootstrapping is called Levelled Fully Homomorphic Encryption (LFHE). In LFHE, the encryption parameters $(N,Q)$ are chosen to provide the desired security and sufficient levels for a circuit with known multiplicative depth.
2.2 Challenges of Using Polynomial Activation
ReLU, defined as $ReLU(x)=\max(0,x)$, is one of the most widely used activation functions in deep learning due to its simplicity and computational efficiency. Its piecewise linear form provides the non-linearity required for neural networks to model complex patterns. However, in FHE settings, ReLU is difficult to implement because it relies on conditional branching, whereas FHE natively supports only additions and multiplications. As a result, any FHE-compatible ReLU must be a polynomial approximation. Higher-degree polynomials produce more accurate approximations but incur larger multiplicative depth. For a degree-$d$ polynomial, the multiplicative depth is $\delta=\left\lceil\log_{2}d\right\rceil+1$, where the $(+1)$ accounts for coefficient multiplication. The lowest multiplicative depth for a ReLU approximation that generalizes to deep CNNs is achieved by PILLAR [11], which uses a degree-4 polynomial $\sum_{i=0}^{4}c_{i}x^{i}$ with $\delta=3$. The theoretical minimum multiplicative depth for an activation function is one, attainable with a quadratic polynomial $x^{2}+c_{1}x+c_{0}$. Such polynomials have been successfully used in shallow CNNs like CryptoNets [13] and LoLa [6], but they fail in deeper models due to the escaping activation problem [12]. Moreover, polynomial approximations are accurate only over bounded input intervals. When used as direct substitutes for ReLU during training, they often cause a substantial drop in model accuracy [12, 15, 11]. This section focuses on two main reasons for this degradation: escaping activations, where intermediate layer outputs leave the polynomial’s approximation range, and coefficient truncation, arising from representing polynomial coefficients in fixed-point format to comply with FHE arithmetic.
Escaping Activations. Let $p_{d}(x)=\sum_{k=0}^{d}a_{k}x^{k}$ be a degree-$d$ polynomial approximation of ReLU, with coefficients $a_{k}\in\mathbb{R}$, $\forall k\in\{0,\dots,d\}$. These coefficients are typically obtained by minimizing the least-squares error between $p_{d}(x)$ and ReLU on a finite set of real-valued points $\{x_{i}\}_{i=1}^{N}\subset[-c,c]$, for some parameter $c>0$: $\min_{a_{0},\ldots,a_{d}}\;\sum_{i=1}^{N}\left(ReLU(x_{i})\;-\;p_{d}(x_{i})\right)^{2}.$ The resulting polynomial approximates ReLU well on [c,c][-c,c] but can deviate substantially outside this range. ReLU exhibits piecewise linear growth:
whereas any polynomial $p_{d}(x)$ with $d\geq 2$ grows as:
$p_{d}(x)\;=\;a_{d}x^{d}\;+\;a_{d-1}x^{d-1}\;+\;\cdots\;+\;a_{1}x\;+\;a_{0}\;=\;\mathcal{O}(x^{d})$
This mismatch in growth rates leads to the escaping activations phenomenon [12]: as inputs propagate through the network, intermediate values can move outside [c,c][-c,c], into regions where $p_{d}(x)$ no longer resembles ReLU. The resulting excessively large activations can cause exploding weights and rapid degradation of model performance, unless the training procedure is specifically adapted to control them.
Coefficient Truncation. In privacy-preserving FHE inference, all computations use fixed-point arithmetic with limited precision. This restriction limits both the range and resolution of representable values, constraining the domain of activation function inputs as well as the magnitudes of their polynomial approximation coefficients. The coefficients $\{a_{k}\}_{k=0}^{d}$ are usually obtained via floating-point least-squares fitting and then quantized for fixed-point evaluation under FHE: $\tilde{a}_{k}\;=\;\frac{\lfloor a_{k}\cdot 2^{b}\rceil}{2^{b}},$ where $b\in\mathbb{N}$ is the number of fractional bits in the fixed-point format and $\lfloor\cdot\rceil$ denotes rounding to the nearest integer. Least-squares fitting often produces small-magnitude coefficients, especially for higher-degree monomials. If $\left|a_{k}\right|<2^{-(b+1)}$, the quantized coefficient $\tilde{a}_{k}$ becomes zero, effectively removing the monomial $x^{k}$ from the approximation. This coefficient truncation [11] changes the polynomial’s shape and can cause substantial deviation from the intended ReLU behavior, even within the approximation interval [c,c][-c,c].
3 Training with Polynomial Approximation
@@ 3.1 Problem Setting
Dataset Configuration. Consider a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{M}$ for a multi-class classification task, where $x_{i}\in\mathbb{R}^{n}$ denotes the feature vector and $y_{i}\in\{1,\dots,K\}$ the corresponding label. The dataset contains $M$ samples distributed across $K$ classes.
ReLU Network. Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{K}$ denote an $L$-layer feed-forward neural network with ReLU activations. The network is parameterized by weight matrices $\{W^{(l)}\}_{l=1}^{L}$, where each $W^{(l)}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ and $n_{l}$ denotes the number of neurons in layer $l$. The network output is defined as $f(x)=h^{(L)}(x)$, where $h^{(0)}(x)=x$ and $h^{(l)}(x)=ReLU(W^{(l)}h^{(l-1)}(x))$ for all $1\leq l\leq L$. The activation function $ReLU(z)$ is applied coordinate-wise as $ReLU(z)=\max(0,z)$.
Polynomial Network. We define a network $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{K}$ with the same architecture as $f$, but replacing each ReLU activation with a degree-$d$ polynomial $p_{d}(z)=\sum_{k=0}^{d}a_{k}z^{k}$ that approximates the ReLU function on a bounded interval [c,c][-c,c]. The polynomial coefficients $\{a_{k}\}_{k=0}^{d}$ are chosen so that the maximum approximation error satisfies: $\max_{z\in[-c,c]}|ReLU(z)-p_{d}(z)|\leq\varepsilon$, where $\varepsilon>0$ is a prescribed approximation tolerance. The resulting network output is given by $g(x)=h^{(L)}_{p_{d}}(x)$, where $h^{(0)}_{p_{d}}(x)=x$ and $h^{(l)}_{p_{d}}(x)=p_{d}(W^{(l)}h^{(l-1)}_{p_{d}}(x))$ for all $1\leq l\leq L$.
Challenges. While polynomial activations enable compatibility with FHE schemes, they introduce several non-trivial optimization challenges, most notably escaping activations and coefficient truncation, as discussed in §2.2. Addressing these issues typically requires careful selection of the polynomial degree $d$, the approximation interval [c,c][-c,c], and coefficient bounds $b$, together with regularization strategies designed to ensure that intermediate activations remain within the valid approximation range throughout training.